OpenStudy (anonymous):

to solve the equation 2sin(2x) = cosx , you should rewite it as ___ a. 2sinx + cosx = 0 b cosx(4sinx-1)= 0 c. cosx(2sinx-1) = 0

OpenStudy (dumbcow):

hint: \[\sin(2x) = 2\sin x \cos x\]

OpenStudy (anonymous):

.......um

OpenStudy (dumbcow):

...plug it in \[4\sin x cosx = \cos x\] ..get right side to be 0

OpenStudy (anonymous):

2 sin 2x = cos x 2(2 sin x cos x) - cos x cos x(4 sin x - 1) = 0 B i think...

OpenStudy (anonymous):

actually is def b